Mary invested 10000 part at 4% and part at 5% for 1 year. How much was invested in each acct if interest earned in the 5% account was 284 greater than in 4% account?
Answer by Boreal(15235)

x invested at 4%
10000-x invested at 5%
.04 x+284=(0.05)(10000-x)
0.04x +284=500-0.05x
0.09x=216
x=2400
10000-x=7600
Check 4% of 2400=$96
5% of 7600=$380
The difference is $284.